what is the center and radius of (x-1)^2+(y+2)^2=25?
Answer by dolly(163)

Comparing the given equation (x-1)^2+(y+2)^2=25 with the standard equation
(x-g)^2+(y-f)^2 = r^2,
we have the centre (g,f) as (1,-2) and the radius r = 5 units